# Direct loading of a Sr magneto-optical trap from a thermal atomic beam

### v2

@@ I Introduction
Direct loading of a grating MOT using a Sr dispenser without employing a Zeeman slower has been demonstrated [A.Sitaram2020], but maintaining UHV conditions requires a differential pumping tube and the number of trapped $^{88}\mathrm{Sr}$ atoms is limited to approximately $10^{6}$. Direct loading of a grating MOT from an Sr atomic oven has also been demonstrated [S.Bondza2022]; however, it requires a slowing laser beam and employs two pumps (an ion pump and a non-evaporable getter pump), and the number of trapped $^{88}\mathrm{Sr}$ atoms is also limited to about $10^{6}$.
@@ II Theory
We first discuss the MOT lifetime determined by collisions with background gas. Under UHV conditions, the background gas is predominantly composed of hydrogen ($\mathrm{H_{2}}$). The loss rate of the number of the trapped atoms due to collisions with $\mathrm{H_{2}}$ is calculated based on the $C_{6}$ coefficient to be $5\times 10^{7}\,\mathrm{s^{-1}\,Torr^{-1}}$ based on the model in Ref. [T.Arpornthip2012] (see Appendix A). Therefore, obtaining an MOT lifetime longer than $1\,\mathrm{s}$ requires a background gas pressure lower than $2\times 10^{-8}\,\mathrm{Torr}$.
@@
Next, we discuss the MOT loading rate. Following the method described in Ref. [M.Barbiero2020], we estimated the capture velocity by simulation using our experimental parameters (see Sec. III), obtaining a value of $\sim 50\,\mathrm{m\,s^{-1}}$. Meanwhile, the atomic beam flux from an oven can be derived from the geometry of capillary tube and the empirical expression for the vapor pressure (see Appendix B). From these considerations, the MOT loading rate calculated for our vacuum system configuration is shown in Fig. 1 (see Appendix C). The figure indicates that even if the capture velocity is $30\,\mathrm{m\,s^{-1}}$, an oven temperature of $\sim 400\,\mathrm{{}^{\circ}C}$ is sufficient to achieve a MOT loading rate of $10^{7}\,\mathrm{atoms\,s^{-1}}$.
@@ III Experimental setup
We first describe the dimensions and features of the vacuum apparatus. As shown in Fig. 4(a), the vacuum chamber is composed of two ICF70 six-way crosses, one of which is equipped with a glass cell ($30\,\mathrm{mm}\times 30\,\mathrm{mm}\times 100\,\mathrm{mm}$), an ion gauge, and an ion pump and the other with an oven and an atomic beam shutter. A key advantage of this system is that UHV conditions are achieved using only a single ion pump (Varian VacIon Plus 55, pumping speed $55\,\mathrm{L\,s^{-1}}$), without the need for a differential pumping tube. The vacuum chamber fits within the size of $500\,\mathrm{mm}\times 500\,\mathrm{mm}\times 150\,\mathrm{mm}$.
The atomic beam is generated by heating a compact oven, as schematically illustrated in Fig. 2, following the design reported in Ref. [M.Schioppo2012]. The beam is collimated using capillary tubes (made of SUS304) with an inner diameter of $0.3\,\mathrm{mm}$, an outer diameter of $0.4\,\mathrm{mm}$, and a length of $10\,\mathrm{mm}$, with 130 capillary tubes installed in total. The oven contains $4\,\mathrm{g}$ of Sr, which is installed in the reservoir with the capillary tubes removed under an argon atmosphere. Using Eq. (14) in Appendix B, we estimate the lifetime of the oven to be about 20 years at a temperature of $400\,\mathrm{{}^{\circ}C}$. The temperature of the oven is measured by a thermocouple at the rear side of the reservoir. The oven is heated by a tantalum wire (diameter of $0.3\,\mathrm{mm}$) insulated by alumina tubes. Thermal isolation from the vacuum chamber using alumina bushes and washers and radiation shielding using a stainless-steel cylindrical reflector enable high oven temperatures with reduced power consumption (Fig. 2). For example, an oven temperature of $400\,\mathrm{{}^{\circ}C}$, requires only $16\,\mathrm{W}$ ($1.6\,\mathrm{A},10\,\mathrm{V}$) of heating power [Fig. 3(a)]. Figure 3(b) represents the dependence of the vacuum pressure measured by the ion gauge on the oven temperature. The vacuum pressure remains in the UHV regime ($\lesssim 1\times 10^{-9}\,\mathrm{Torr}$) at temperatures below $380\,\mathrm{{}^{\circ}C}$, presumably owing to the getter effect of metallic Sr deposited on the surfaces of the vacuum chamber and the stainless-steel reflector.
@@ IV Results and Discussion
To estimate the loss rate of trapped atoms due to collisions with background gas, we block the atomic beam with the beam shutter and monitor the decay curve of the the number of trapped atoms. Figure 6 shows the decay curve of the trapped atom number at an oven temperature of $325\,\mathrm{{}^{\circ}C}$. The initial fast decay is due to two-body collisions, which is followed by the slow exponential decay caused by collisions with background gas. Figure 7(a) shows the dependence of the loss rate due to collisions with background gas on the oven temperature, obtained by fitting the latter decay with an exponential function. As seen in the figure, this loss rate increases with oven temperature. Figure 7(b) demonstrates a linear relationship between background gas pressure measured by the ion gauge and the loss rate, with a coefficient of $\alpha_{\text{fit}}=1.67(4)\times 10^{8}\,\mathrm{s^{-1}\,Torr^{-1}}$. In contrast, the coefficient calculated using Arpornthip’s expression [T.Arpornthip2012] and the $C_{6}$ coefficients from Ref. [M.Abdel-Hafiz2019] is $\alpha=5.2(5)\times 10^{7}\,\mathrm{s^{-1}\,Torr^{-1}}$ (see Appendix A), giving a ratio of $\alpha_{\text{fit}}/\alpha=3.2(3)$. This indicates that the background gas pressure in the glass cell is approximately three times higher than that measured by the ion gauge. A plausible explanation is a poor conductance ($\sim 5\,\mathrm{L\,s^{-1}}$) of the neck of the glass cell.
@@ Appendix B Atomic beam flux
We determine the atomic beam flux from the oven by measuring the absorption profile of the atomic beam emitted from the oven as shown in Fig. 4. The arguments presented in this section are primarily based on Ref. [M.Schioppo2012].
@@
where $n$ is the atomic beam density, $r_{i}$ is the isotopic abundance of isotope $i$ for bosons ($^{84}\mathrm{Sr},{}^{86}\mathrm{Sr},{}^{88}\mathrm{Sr}$) and the relative contribution of each hyperfine line $i$ for fermion ($^{87}\mathrm{Sr}$), $\sigma_{0}$ is the resonant scattering cross section without Doppler broadening $\left[\sigma_{0}=3\lambda^{2}/(2\pi)\right]$, $l_{\text{int}}$ is the thickness of the atomic beam along the probe direction (set to $l_{\text{int}}=5\,\mathrm{mm}$ from the aperture size), $\sigma_{t}$ is the transverse Doppler width, $v_{t}$ is the transverse atomic velocity, $\nu_{0}^{i}$ is the resonance frequency of isotope $i$ for bosons and of each hyperfine line $i$ for fermion, $\lambda$ is the probe wavelength ($461\,\mathrm{nm}$), and $\Gamma$ is the natural linewidth (FWHM) of the transition ($2\pi\times 30.2\,\mathrm{MHz}$). Details on isotope shifts and relative transition strength of $^{87}\mathrm{Sr}$ can be found in Refs. [NIST, S.Mauger2007]. The absorption profile of the atomic beam is then fitted using Eqs. (6) and (7), with $n$ and $\sigma_{t}$ as fitting parameters (see Fig. 4). From the fitting results, the atomic beam density $n$ is obtained.
@@
where $\Omega_{\text{aperture}}$ is the solid angle subtended by the aperture as seen from the oven. For small divergence angles it can be approximated by
@@
where $W=(8a/3L)/(1+8a/3L)$ the capillary length is the transmission probability (Clausing factor) of the capillaries, with $L=10\,\mathrm{mm}$ the capillary length.
@@ Appendix D Dependence of the trapped atom number on the loading rate
According to Ref. [T.Dinneen1999], in constant-volume regime, the density-normalized two-body loss coefficient $\beta$ ($\mathrm{cm^{3}\,s^{-1}}$) is related to $\tilde{\beta}$ as

### v3

@@ I Introduction
Direct loading of a grating MOT using a Sr dispenser without employing a Zeeman slower has been demonstrated [A.Sitaram2020], but maintaining UHV conditions requires a differential pumping tube and the number of trapped $^{88}\mathrm{Sr}$ atoms is limited to approximately $10^{6}$. Direct loading of a grating MOT from a Sr atomic oven has also been demonstrated [S.Bondza2022]; however, it requires a slowing laser beam and employs two pumps (an ion pump and a non-evaporable getter pump), and the number of trapped $^{88}\mathrm{Sr}$ atoms is also limited to about $10^{6}$.
@@ II Theory
We first discuss the MOT lifetime determined by collisions with background gas. Under UHV conditions, the background gas is predominantly composed of hydrogen ($\mathrm{H_{2}}$). The loss rate of the number of the trapped atoms due to collisions with $\mathrm{H_{2}}$ is calculated to be $5\times 10^{7}\,\mathrm{s^{-1}\,Torr^{-1}}$ based on the model in Ref. [T.Arpornthip2012] (see Appendix A). Therefore, obtaining an MOT lifetime longer than $1\,\mathrm{s}$ requires a background gas pressure lower than $2\times 10^{-8}\,\mathrm{Torr}$.
@@
Next, we discuss the MOT loading rate. Following the method described in Ref. [M.Barbiero2020], we estimated the capture velocity by simulation using our experimental parameters (see Sec. III), obtaining a value of $\sim 50\,\mathrm{m\,s^{-1}}$. Meanwhile, the atomic beam flux from an oven can be derived from the geometry of the capillary tubes and the empirical expression for the vapor pressure (see Appendix B). From these considerations, the MOT loading rate calculated for our vacuum system configuration is shown in Fig. 1 (see Appendix C). The figure indicates that even if the capture velocity is $30\,\mathrm{m\,s^{-1}}$, an oven temperature of $\sim 400\,\mathrm{{}^{\circ}C}$ is sufficient to achieve a MOT loading rate of $10^{7}\,\mathrm{atoms\,s^{-1}}$.
@@ III Experimental setup
We first describe the dimensions and features of the vacuum apparatus. As shown in Fig. 4(a), the vacuum chamber is composed of two ICF70 six-way crosses, one of which is equipped with a glass cell ($30\,\mathrm{mm}\times 30\,\mathrm{mm}\times 100\,\mathrm{mm}$), an ion gauge, and an ion pump and the other with an oven and an atomic beam shutter. A key advantage of this system is that UHV conditions are achieved using only a single ion pump (Varian VacIon Plus 55, pumping speed $55\,\mathrm{L\,s^{-1}}$), without the need for a differential pumping tube. The vacuum chamber fits within a volume of $500\,\mathrm{mm}\times 500\,\mathrm{mm}\times 150\,\mathrm{mm}$.
The atomic beam is generated by heating a compact oven, as schematically illustrated in Fig. 2, following the design reported in Ref. [M.Schioppo2012]. The beam is collimated using capillary tubes (made of SUS304) with an inner diameter of $0.3\,\mathrm{mm}$, an outer diameter of $0.4\,\mathrm{mm}$, and a length of $10\,\mathrm{mm}$, with 130 capillary tubes installed in total. The oven contains $4\,\mathrm{g}$ of Sr, which is installed in the reservoir with the capillary tubes removed under an argon atmosphere. Using Eq. (14) in Appendix B, we estimate the lifetime of the oven to be about 20 years at a temperature of $400\,\mathrm{{}^{\circ}C}$. The temperature of the oven is measured by a thermocouple at the rear side of the reservoir. The oven is heated by a tantalum wire (diameter of $0.3\,\mathrm{mm}$) insulated by alumina tubes. Thermal isolation from the vacuum chamber using alumina bushes and washers and radiation shielding using a stainless-steel cylindrical reflector enable high oven temperatures with reduced power consumption (Fig. 2). For example, an oven temperature of $400\,\mathrm{{}^{\circ}C}$ requires only $16\,\mathrm{W}$ ($1.6\,\mathrm{A},10\,\mathrm{V}$) of heating power [Fig. 3(a)]. Figure 3(b) represents the dependence of the vacuum pressure measured by the ion gauge on the oven temperature. The vacuum pressure remains in the UHV regime ($\lesssim 1\times 10^{-9}\,\mathrm{Torr}$) at temperatures below $380\,\mathrm{{}^{\circ}C}$, presumably owing to the getter effect of metallic Sr deposited on the surfaces of the vacuum chamber and the stainless-steel reflector.
@@ IV Results and Discussion
To estimate the loss rate of trapped atoms due to collisions with background gas, we block the atomic beam with the beam shutter and monitor the decay curve of the number of trapped atoms. Figure 6 shows the decay curve of the trapped atom number at an oven temperature of $325\,\mathrm{{}^{\circ}C}$. The initial fast decay is due to two-body collisions, which is followed by the slow exponential decay caused by collisions with background gas. Figure 7(a) shows the dependence of the loss rate due to collisions with background gas on the oven temperature, obtained by fitting the latter decay with an exponential function. As seen in the figure, this loss rate increases with oven temperature. Figure 7(b) demonstrates a linear relationship between background gas pressure measured by the ion gauge and the loss rate, with a coefficient of $\alpha_{\text{fit}}=1.67(4)\times 10^{8}\,\mathrm{s^{-1}\,Torr^{-1}}$. In contrast, the coefficient calculated using Arpornthip’s expression [T.Arpornthip2012] and the $C_{6}$ coefficients from Ref. [M.Abdel-Hafiz2019] is $\alpha=5.2(5)\times 10^{7}\,\mathrm{s^{-1}\,Torr^{-1}}$ (see Appendix A), giving a ratio of $\alpha_{\text{fit}}/\alpha=3.2(3)$. This indicates that the background gas pressure in the glass cell is approximately three times higher than that measured by the ion gauge. A plausible explanation is a poor conductance ($\sim 5\,\mathrm{L\,s^{-1}}$) of the neck of the glass cell.
@@ Appendix B Atomic beam flux
We determine the atomic beam flux by measuring the absorption profile of the atomic beam emitted from the oven as shown in Fig. 4. The arguments presented in this section are primarily based on Ref. [M.Schioppo2012].
@@
where $n$ is the atomic beam density, $r_{i}$ is the isotopic abundance of isotope $i$ for bosons ($^{84}\mathrm{Sr},{}^{86}\mathrm{Sr},{}^{88}\mathrm{Sr}$) and the relative contribution of each hyperfine line $i$ for fermion ($^{87}\mathrm{Sr}$), $\sigma_{0}=3\lambda^{2}/(2\pi)$ is the resonant scattering cross section without Doppler broadening, $l_{\text{int}}$ is the thickness of the atomic beam along the probe direction (set to $l_{\text{int}}=5\,\mathrm{mm}$ from the aperture size), $\sigma_{t}$ is the transverse Doppler width, $v_{t}$ is the transverse atomic velocity, $\nu_{0}^{i}$ is the resonance frequency of isotope $i$ for bosons and of each hyperfine line $i$ for fermion, $\lambda$ is the probe wavelength ($461\,\mathrm{nm}$), and $\Gamma$ is the natural linewidth (FWHM) of the transition ($2\pi\times 30.2\,\mathrm{MHz}$). Details on isotope shifts and relative transition strengths of $^{87}\mathrm{Sr}$ can be found in Ref. [S.Mauger2007]. The absorption profile of the atomic beam is then fitted using Eqs. (6) and (7), with $n$ and $\sigma_{t}$ as fitting parameters (see Fig. 4). From the fitting results, the atomic beam density $n$ is obtained.
@@
where $\Omega_{\text{aperture}}$ is the solid angle subtended by the aperture as seen from the oven. For small divergence angles, it can be approximated by
@@
where $W=(8a/3L)/(1+8a/3L)\sim 0.04$ is the transmission probability (Clausing factor) of the capillaries, with $L=10\,\mathrm{mm}$ being the capillary length.
@@ Appendix D Dependence of the trapped atom number on the loading rate
According to Ref. [T.Dinneen1999], in the constant-volume regime, the density-normalized two-body loss coefficient $\beta$ ($\mathrm{cm^{3}\,s^{-1}}$) is related to $\tilde{\beta}$ as